OpenStudy (anonymous):

can somebody please help me simplify 8 -1/3 ? these are the choices.. -1/2 1/2 2 –2

OpenStudy (lalaly):

are u sure the options are right?

OpenStudy (akshay_budhkar):

none of these i believe

OpenStudy (anonymous):

agree

OpenStudy (anonymous):

do you mean 8 to the power 1/3?

OpenStudy (akshay_budhkar):

23/3 is your answer

OpenStudy (anonymous):

to the -1/3?

OpenStudy (akshay_budhkar):

oh ya maybe jimmy is right

OpenStudy (akshay_budhkar):

didnt think of it

OpenStudy (anonymous):

yeah im sorry i meant 8 to the power of -1/3

OpenStudy (lalaly):

1/2

OpenStudy (akshay_budhkar):

1/2 then

OpenStudy (anonymous):

in which case it is 1/2

OpenStudy (akshay_budhkar):

jimmy deserves to be a master hatts off

OpenStudy (anonymous):

1/2

OpenStudy (anonymous):

lol ty akshay!

OpenStudy (anonymous):

\[8^{\frac{1}{3}}=\sqrt[3]{8}=2\] so \[8^{-\frac{1}{3}}=\frac{1}{2}\] because the denominator means the root, and the minus sign means take the reciprocal